# Strict area law entanglement versus chirality

### v2

@@ 2.  Preliminaries
where $(S_{X})_{\left|\Psi\right\rangle}\equiv S(\rho_{X})$. With Eq. (2.1) one can check that a strict area law state satisfies bulk A1. On the other hand, A1 states include more than strict area law states, since A1 could still hold if there are additional contributions to Eq. (2.1)  [^2] In general, bulk A1 allows any additional short-range entanglement contributions in the form of sum of local terms along $\partial A$ as discussed in [23]. An explicit state satisfying A1 but not strict area law can be obtained by starting with a state with strict area law and tensoring in EPR pairs $\sqrt{p_{uv}}\left|0_{u}0_{v}\right\rangle+\sqrt{1-p_{uv}}\left|1_{u}1_{v}\right\rangle$ of varying strength $p_{uv}$ along each link $<uv>$ of the lattice. The whole state satisfies A1 still, but the area law coefficient will vary in space according to the choices of $p$. Moreover, sharp corners of the region can make additional contributions to Eq. (2.1), which are consistent with A1. In fact, we show in [24] that such a corner contribution is a necessary consequence of $c_{-}\neq 0$. However, as a consequence of the no-go theorem 3.1, this violation of the exact area law cannot be the only one that results from $c_{-}\neq 0$ – since a corner contribution is consistent with A1, further violation is required.. An important fact about a state that satisfies bulk A1 is
@@ 4.  Instantaneous modular flow
$J(A,B,C)_{\left|\Psi\right\rangle}=2{\rm Im\hskip 1.00006pt}\Big{[}\partial_{x}\partial_{y}\mathcal{F}^{A,B,C}_{J}(x,y)_{\left|\Psi\right\rangle}\Big{]}_{x,y=0},$ (4.4)
where $\mathcal{F}^{A,B,C}_{J}(x,y)_{\left|\Psi\right\rangle}\equiv\left(\mathcal{I}_{AB}(x)\left|\Psi\right\rangle,\mathcal{I}_{BC}(y)\left|\Psi\right\rangle\right)$ [^7] Here we use (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) to denote state overlap: $(\left|\varphi\right\rangle,\left|\psi\right\rangle)\equiv\langle\varphi|\psi\rangle$. Similarly, one can compute $\Sigma(A,B,C)_{\left|\Psi\right\rangle}$ via
@@
with $\mathcal{F}^{A,B,C}_{\Sigma}(x,y)_{\left|\Psi\right\rangle}\equiv\left(\mathcal{I}_{AB}(x)\left|\Psi\right\rangle,U_{BC}(y)\left|\Psi\right\rangle\right)$. This way of writing $J$ or $\Sigma$ is less singular in that it avoids using modular Hamiltonians; furthermore, the more general quantity ${\cal F}$ turns out to be interesting in itself and more tractable than $J$ or $\Sigma$.
@@ 5.  Proofs of the no-go theorems
The proof of Theorem 3.1 is by contradiction. If there were a state $\left|\Psi\right\rangle$ with the properties described in Theorem 3.1, then based on $\left|\Psi\right\rangle$ one could construct a family of states $\{\rho_{X}(t)\}_{t\in\mathbb{R}}$ on a local Hilbert space $\mathcal{H}_{X}=\otimes_{v\in X}\mathcal{H}_{v}$ of a region $X$ (a collection of finite number of sites in the lattice), such that the entanglement entropies of these states $S(\rho_{X}(t))$ would be unbounded as a function of $t\in\mathbb{R}$. This contradicts the setup that $\mathcal{H}_{X}$ has a finite dimension $d_{X}$, because entanglement entropies of any state within $\mathcal{H}_{X}$ take values in range $[0,\ln d_{X}]$.
@@
Therefore, $S_{BC}(t)$ is unbounded as a function of $t\in\mathbb{R}$ from both sides. This is a contradiction since $\rho_{BC}(t)$ in $\mathcal{H}_{BC}$ of dimension $d_{BC}$ should have entanglement entropies within $[0,\ln d_{BC}]$.
@@ 6.  Discussion
Our no-go theorems reveal that chirality, in the form of $c_{-}\neq 0$ or $\sigma_{xy}\neq 0$, constrains the local entanglement properties of a representative state of a chiral gapped phase [^10] We comment that one might hope to combine the recent result of [33], which says that any state satisfying A0 and A1 has a commuting projector parent Hamiltonian, with the Hamiltonian no-go theorems of [6, 7, 8], to provide a proof of a related result.. Loosely speaking, our no-go theorems indicate that it is impossible to construct a state in a Hilbert space with finite local dimensions (or alternatively, with finite local entanglement entropies) that represent a zero correlation length fixed point of a chiral gapped phase. [^11] This also aligns with various difficulties when people attempt to construct a chiral gapped state, such as using PEPS [Hasik:2022, Wahl:2013rha] or MERA [Chu:2023ven]. We leave the detailed connections to be investigated in the future. Moreover, the no-go theorems and their proofs might give a hint on studying the subleading correction terms to the area law entanglement formula Eq. (2.1). For a representative of a chiral gapped phase in a finite-dimensional local Hilbert space, we expect that the bulk A1 is approximately satisfied with violations decreasing as the subsystem sizes are enlarged, consistent with the decay of the subleading corrections to Eq. (2.1). On such a state, it is expected that $J$ and $\Sigma$ compute $c_{-}$ and $\sigma_{xy}$ also with decreasing error for larger subsystem sizes. One example of such a case is studied in the numerical section of [22]. In fact, the techniques used in the proofs might be used to relate the violation of the Markov flow equation Eq. (4.6), which is a signature of the violation of bulk A1, to the deviations of $J$ and $\Sigma$ from the expected values. The detailed relation is left for future investigation.
@@ C.2.1 Basic facts about symmetry charge operators
where the $\Rightarrow$ can be seen by tracing out $\overline{X}$. Then one can see $U_{X}(t)K_{X}U_{X}(t)^{\dagger}=K_{X}$ since $K_{X}=-\ln\rho_{X}$. Taking the first derivative of $t$ at $t=0$, we obtain $[K_{X},Q_{X}]=0$. With $Q_{AB}=Q_{A}+Q_{B}$, we can see $[K_{A},Q_{AB}]=0$. To see $[K_{A},Q_{AB}^{2}]=0$, we can utilize the Fact (1) to write $Q_{AB}=Q_{A}+Q_{B}$. Therefore,
@@
The three commutators above all vanish because $[K_{A},Q_{A}]=0$ and $[K_{A},Q_{B}]=0$.
@@ C.2.2 Properties of $\Sigma(A,B,C)$
In the derivation, we first utilize Eq. (C.11) to replace $K_{AB}\left|\Psi\right\rangle$. Then, in the second line to the third line, we flip the support of $K_{ABC},K_{BC}$ to their compliments, which have no overlap with $Q_{B}Q_{C}$ and hence the first two commutators in the third line vanish. To show the remaining term vanishes $\langle[K_{B},Q_{B}Q_{C}]\rangle=0$, we utilize $[K_{B},Q_{B}]=0$ [Fact C.3 (2)] in the second last line. ∎
@@
    because $[K_{AB},Q_{A}Q_{z^{\prime}}]+[K_{AB},Q_{B}Q_{z^{\prime}}]=[K_{AB},Q_{AB}Q_{z^{\prime}}]=0$. Then, notice $I(z^{\prime}:B|A)=0$, which enable us to replace $K_{AB}$ in the commutator using
@@
    In the equation above, $\langle[K_{Az^{\prime}},Q_{A}Q_{z^{\prime}}]\rangle=\langle[K_{ABz^{\prime}},Q_{A}Q_{z^{\prime}}]\rangle=0$ because one can flip the support of $K_{Az^{\prime}}$ and $K_{ABz^{\prime}}$ to their compliments so that they have no overlap with $Az^{\prime}$. $\langle[K_{A},Q_{A}Q_{z^{\prime}}]\rangle=0$ is simply because $[K_{A},Q_{A}Q_{z^{\prime}}]=[K_{A},Q_{A}]Q_{z^{\prime}}=0$ with [Fact C.3 (2)].
@@
Now we shall show the second term in the right-hand of Eq. (E.11) vanishes. As a starter, we emphasize that as a consequence of Lemma E.1 $\left|\Psi^{\prime}(t)\right\rangle$ is also symmetric under the same $\textsf{U}(1)$ operator generated by $Q=\sum_{v}Q_{v}$ as the one for $\left|\Psi\right\rangle$, therefore both Fact C.3 and Lemma C.4 are applicable for $\left|\Psi^{\prime}(t)\right\rangle$. Firstly, we can use Fact C.3 (2) to conclude $[K^{\prime}_{AB},Q_{AB}Q_{M}]=0$. With $Q_{B}=Q_{AB}-Q_{A}$ and $[K^{\prime}_{AB},Q_{AB}]=0$, we can write the second term in Eq. (E.11) as
@@ Appendix F Generalizations of the no-go theorems
The no-go theorems 3.1 and 3.2 both involve the finite local Hilbert space dimensions condition. This condition can be relaxed into the condition that $\left|\Psi\right\rangle$ has finite local entanglement entropies and finite local charge fluctuations (i.e. $(S_{A})_{\left|\Psi\right\rangle}<\infty$ and $\langle\Psi|Q_{A}^{2}|\Psi\rangle-\langle\Psi|Q_{A}|\Psi\rangle^{2}<\infty$ for any local region $A$). [^14] We left it a question on if the finiteness of entropy implies the finiteness of charge fluctuation.
@@
  • Lemma 4.2 is still valid and therefore, one can still do Markov decomposition move. This is because if $\left|\Psi\right\rangle$ has finite local entanglement entropies, even if it is in an infinite dimensional Hilbert space, the structure theorem Eq. (B.2) of a Markov state $\rho_{ABC}$ from $\left|\Psi\right\rangle$ still holds [Jen2006], and proof of Lemma 4.2 given in App. B is still applicable.
@@
Since $(S_{BC})_{\left|\Psi(t)\right\rangle}\geq 0,\forall t\in\mathbb{R}$ and one can take $\pi c_{-}t/3\to-\infty$ by letting $t\to-\infty$ (or +$\infty$+ ∞) if $c_{-}>0$ (or $<0$), $(S_{BC})_{\left|\Psi\right\rangle}$ must be infinite, which contradicts to the finite local entropies assumption of $\left|\Psi\right\rangle$.

### v3

@@ 2.  Preliminaries
where $(S_{X})_{\left|\Psi\right\rangle}\equiv S(\rho_{X})$. With Eq. (2.1) one can check that a strict area law state satisfies bulk A1. On the other hand, A1 states include more than strict area law states, since A1 could still hold if there are additional contributions to Eq. (2.1)  [^2] In general, bulk A1 allows any additional short-range entanglement contributions in the form of sum of local terms along $\partial A$ as discussed in [23]. An explicit state satisfying A1 but not strict area law can be obtained by starting with a state with strict area law and tensoring in EPR pairs $\sqrt{p_{uv}}\left|0_{u}0_{v}\right\rangle+\sqrt{1-p_{uv}}\left|1_{u}1_{v}\right\rangle$ of varying strength $p_{uv}$ along each link <uv><uv> of the lattice. The whole state satisfies A1 still, but the area law coefficient will vary in space according to the choices of $p$. Moreover, sharp corners of the region can make additional contributions to Eq. (2.1), which are consistent with A1. In fact, we show in [24] that such a corner contribution is a necessary consequence of $c_{-}\neq 0$. However, as a consequence of the no-go theorem 3.1, this violation of the exact area law cannot be the only one that results from $c_{-}\neq 0$ – since a corner contribution is consistent with A1, further violation is required.. An important fact about a state that satisfies bulk A1 is
@@ 4.  Instantaneous modular flow
$J(A,B,C)_{\left|\Psi\right\rangle}=2{\rm Im\thinspace}\Big[\partial_{x}\partial_{y}\mathcal{F}^{A,B,C}_{J}(x,y)_{\left|\Psi\right\rangle}\Big]_{x,y=0},$ (4.4)
where $\mathcal{F}^{A,B,C}_{J}(x,y)_{\left|\Psi\right\rangle}\equiv\left(\mathcal{I}_{AB}(x)\left|\Psi\right\rangle,\mathcal{I}_{BC}(y)\left|\Psi\right\rangle\right)$ [^7] Here we use $(\cdot,\cdot)$ to denote state overlap: $(\left|\varphi\right\rangle,\left|\psi\right\rangle)\equiv\langle\varphi|\psi\rangle$. Similarly, one can compute $\Sigma(A,B,C)_{\left|\Psi\right\rangle}$ via
@@
[…]
@@ 5.  Proofs of the no-go theorems
The proof of Theorem 3.1 is by contradiction. If there were a state $\left|\Psi\right\rangle$ with the properties described in Theorem 3.1, then based on $\left|\Psi\right\rangle$ one could construct a family of states $\{\rho_{X}(t)\}_{t\in\mathbb{R}}$ on a local Hilbert space $\mathcal{H}_{X}=\otimes_{v\in X}\mathcal{H}_{v}$ of a region $X$ (a collection of finite number of sites in the lattice), such that the entanglement entropies of these states $S(\rho_{X}(t))$ would be unbounded as a function of $t\in\mathbb{R}$. This contradicts the setup that $\mathcal{H}_{X}$ has a finite dimension $d_{X}$, because entanglement entropies of any state within $\mathcal{H}_{X}$ take values in range $0,\ln d_{X}]$.
@@
Therefore, $S_{BC}(t)$ is unbounded as a function of $t\in\mathbb{R}$ from both sides. This is a contradiction since $\rho_{BC}(t)$ in $\mathcal{H}_{BC}$ of dimension $d_{BC}$ should have entanglement entropies within $0,\ln d_{BC}]$.
@@ 6.  Discussion
Our no-go theorems reveal that chirality, in the form of $c_{-}\neq 0$ or $\sigma_{xy}\neq 0$, constrains the local entanglement properties of a representative state of a chiral gapped phase [^10] We comment that one might hope to combine the recent result of [33], which says that any state satisfying A0 and A1 has a commuting projector parent Hamiltonian, with the Hamiltonian no-go theorems of [6, 7, 8], to provide a proof of a related result.. Loosely speaking, our no-go theorems indicate that it is impossible to construct a state in a Hilbert space with finite local dimensions (or alternatively, with finite local entanglement entropies) that represent a zero correlation length fixed point of a chiral gapped phase. [^11] This also aligns with various difficulties when people attempt to construct a chiral gapped state, such as using PEPS [34, 35] or MERA [36]. We leave the detailed connections to be investigated in the future. Moreover, the no-go theorems and their proofs might give a hint on studying the subleading correction terms to the area law entanglement formula Eq. (2.1). For a representative of a chiral gapped phase in a finite-dimensional local Hilbert space, we expect that the bulk A1 is approximately satisfied with violations decreasing as the subsystem sizes are enlarged, consistent with the decay of the subleading corrections to Eq. (2.1). On such a state, it is expected that $J$ and $\Sigma$ compute $c_{-}$ and $\sigma_{xy}$ also with decreasing error for larger subsystem sizes. One example of such a case is studied in the numerical section of [22]. In fact, the techniques used in the proofs might be used to relate the violation of the Markov flow equation Eq. (4.6), which is a signature of the violation of bulk A1, to the deviations of $J$ and $\Sigma$ from the expected values. The detailed relation is left for future investigation.
@@ C.2.1 Basic facts about symmetry charge operators
where the $\Rightarrow$ can be seen by tracing out $\overline{X}$. Then one can see $U_{X}(t)K_{X}U_{X}(t)^{\dagger}=K_{X}$ since $K_{X}=-\ln\rho_{X}$. Taking the first derivative of $t$ at $t=0$, we obtain $K_{X},Q_{X}]=0$. With $Q_{AB}=Q_{A}+Q_{B}$, we can see $K_{A},Q_{AB}]=0$. To see $K_{A},Q_{AB}^{2}]=0$, we can utilize the Fact (1) to write $Q_{AB}=Q_{A}+Q_{B}$. Therefore,
@@
The three commutators above all vanish because $K_{A},Q_{A}]=0$ and $K_{A},Q_{B}]=0$.
@@ C.2.2 Properties of $\Sigma(A,B,C)$
In the derivation, we first utilize Eq. (C.11) to replace $K_{AB}\left|\Psi\right\rangle$. Then, in the second line to the third line, we flip the support of $K_{ABC},K_{BC}$ to their compliments, which have no overlap with $Q_{B}Q_{C}$ and hence the first two commutators in the third line vanish. To show the remaining term vanishes $\langle[K_{B},Q_{B}Q_{C}]\rangle=0$, we utilize $K_{B},Q_{B}]=0 [$Fact C.3 (2)] in the second last line. ∎
@@
    because $K_{AB},Q_{A}Q_{z^{\prime}}]+[K_{AB},Q_{B}Q_{z^{\prime}}]=[K_{AB},Q_{AB}Q_{z^{\prime}}]=0$. Then, notice $I(z^{\prime}:B|A)=0$, which enable us to replace $K_{AB}$ in the commutator using
@@
    In the equation above, $\langle[K_{Az^{\prime}},Q_{A}Q_{z^{\prime}}]\rangle=\langle[K_{ABz^{\prime}},Q_{A}Q_{z^{\prime}}]\rangle=0$ because one can flip the support of $K_{Az^{\prime}}$ and $K_{ABz^{\prime}}$ to their compliments so that they have no overlap with $Az^{\prime}$. $\langle[K_{A},Q_{A}Q_{z^{\prime}}]\rangle=0$ is simply because $K_{A},Q_{A}Q_{z^{\prime}}]=[K_{A},Q_{A}]Q_{z^{\prime}}=0$ with [Fact C.3 (2)].
@@
Now we shall show the second term in the right-hand of Eq. (E.11) vanishes. As a starter, we emphasize that as a consequence of Lemma E.1 $\left|\Psi^{\prime}(t)\right\rangle$ is also symmetric under the same $\textsf{U}(1)$ operator generated by $Q=\sum_{v}Q_{v}$ as the one for $\left|\Psi\right\rangle$, therefore both Fact C.3 and Lemma C.4 are applicable for $\left|\Psi^{\prime}(t)\right\rangle$. Firstly, we can use Fact C.3 (2) to conclude $K^{\prime}_{AB},Q_{AB}Q_{M}]=0$. With $Q_{B}=Q_{AB}-Q_{A}$ and $K^{\prime}_{AB},Q_{AB}]=0$, we can write the second term in Eq. (E.11) as
@@ Appendix F Generalizations of the no-go theorems
The no-go theorems 3.1 and 3.2 both involve the finite local Hilbert space dimensions condition. This condition can be relaxed into the condition that $\left|\Psi\right\rangle$ has finite local entanglement entropies and finite local charge fluctuations (i.e. $(S_{A})_{\left|\Psi\right\rangle}<\infty$ and $\langle\Psi|Q_{A}^{2}|\Psi\rangle-\langle\Psi|Q_{A}|\Psi\rangle^{2}<\infty$ for any local region AA). [^14] We left it a question on if the finiteness of entropy implies the finiteness of charge fluctuation.
@@
  • Lemma 4.2 is still valid and therefore, one can still do Markov decomposition move. This is because if $\left|\Psi\right\rangle$ has finite local entanglement entropies, even if it is in an infinite dimensional Hilbert space, the structure theorem Eq. (B.2) of a Markov state $\rho_{ABC}$ from $\left|\Psi\right\rangle$ still holds [43], and proof of Lemma 4.2 given in App. B is still applicable.
@@
Since $(S_{BC})_{\left|\Psi(t)\right\rangle}\geq 0,\forall t\in\mathbb{R}$ and one can take $\pi c_{-}t/3\to$-$\infty$ by letting $t\to$-$\infty$ (or $\infty$) if $c_{-}>0$ (or <0<0), $(S_{BC})_{\left|\Psi\right\rangle}$ must be infinite, which contradicts to the finite local entropies assumption of $\left|\Psi\right\rangle$.